OpenStudy (anonymous):

Choose if the following function is even, odd or neither. f(x) = x 3 even odd neither

OpenStudy (across):

A function \(f\) is even if \(f(-x)=f(x)\) and odd if \(f(-x)=-f(x)\).

Directrix (directrix):

@david93 Crank out f(x) = (x)^3 for x = -x. Hint: -x would replace x. Consider the result of (-x) ^3 as compared to the result of - f(x) which would be - ( (x)^3). Are they the same?